write the equation of the line that passes through (-2,8) and has a slope of -3/5
Answer by Mathtut(3670)

point slope for y-k=m(x-h) where m is the slope and (h,k) is a point on the line
so our point is (-2,8) y-8=3/5(x+2) ---> y-8=3/5x+6/5----> y=3/5x+46/5
or 5y-3x-46=0